OpenStudy (anonymous):

Solve by simulating the problem. You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly? Type your answer below using complete sentences.

OpenStudy (anonymous):

I can't do this for you.

OpenStudy (anonymous):

well at least Someone can explain probabilities which i really hate :P

OpenStudy (anonymous):

nah bye dude.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

its this (1/4)(1/4)(1/4) =1/16

OpenStudy (anonymous):

thanks @CO_olBoY

OpenStudy (anonymous):

sry i'm wrong its 1/64

OpenStudy (anonymous):

not 1/16

OpenStudy (kropot72):

The probability of guessing the answers correctly to 3 out of the 5 questions is found from the binomial distribution as follows: \[P(3\ correct\ out\ of\ 5)=5C3\times(\frac{1}{4})^{3}\times(\frac{3}{4})^{2}\]

OpenStudy (kropot72):

In words, the required probability is the number of combinations of five questions taken three at a time, multiplied by the probability of guessing 3 questions correctly, multiplied by the probability of guessing 2 questions incorrectly.